OpenStudy (anonymous):

I think this is right but could someone just verify my answer. The question is "find the values of c and d such that the function is continuous on (-infinity, +infinity). The function will be posted below.

OpenStudy (anonymous):

F(x)= 2x if x<1, cx^2 +d if 1</=x</=2, 4x if 2<x

OpenStudy (anonymous):

Btw I got c=2 and d=0 just looking for verification.

OpenStudy (amistre64):

does 2(1)= c(1)^2+d and 4(2) = c(2)^2 + d when c=2 and d=0?

OpenStudy (anonymous):

yes thank you